OpenStudy (anonymous):

I JUST NEED AN ANSWER Let f(x) = x - 3 and g(x) = 2x + 4. What is f(g(x)) ? a; 2x+7 b; 2x-2 c; 2x+1 d; 3x+1

OpenStudy (faiqraees):

subtsitute g(x) into f(x)

OpenStudy (mathmale):

@GmasterJoe: New here? Welcome. Please note that OpenStudy rules forbid both asking for direct answers and giving them out. Let f(x) = x - 3 and g(x) = 2x + 4. What is f(g(x)) ? a; 2x+7 b; 2x-2 c; 2x+1 d; 3x+1 f(g(x)) signifies a "composition." You are to use the function g(x)=2x+4 as the input to function f(x), in that order. If f(x)=x-3, and x=2, then f(2)=2-3=-1. Now, if f(x)= x-3, substitute g(x) for x on the left side and substitute (2x+4) for x on the right side. Show your work, please, and simplify your answer.